# Mahyar Habibi,1 Dirk Hovy,2 Carlo Schwarz1

### v1

@@ Abstract
There is an ongoing debate about how to moderate toxic speech on social media and how content moderation affects online discourse. We propose and validate a methodology for measuring the content-moderation-induced distortions in online discourse using text embeddings from computational linguistics. We test our measure on a representative dataset of 5 million US political Tweets and find that removing toxic Tweets distorts online content. This finding is consistent across different embedding models, toxicity metrics, and samples. Importantly, we demonstrate that content-moderation-induced distortions are not caused by the toxic language. Instead, we show that, as a side effect, content moderation shifts the mean and variance of the embedding space, distorting the topic composition of online content. Finally, we propose an alternative approach to content moderation that uses generative Large Language Models to rephrase toxic Tweets to preserve their salvageable content rather than removing them entirely. We demonstrate that this rephrasing strategy reduces toxicity while minimizing distortions in online content.
@@
Significance Statement
Automated content moderation on social media platforms aims to reduce toxic content but is frequently criticized for distorting online discourse. However, the extent of these distortions is unknown. We present a novel measure to quantify content-moderation-induced distortions in online discourse. Analyzing 5 million US political tweets, we show that content moderation causes significant changes in the conceptual space of online debate, comparable to removing entire topics (e.g., abortion, BLM) from the platform. Finally, we show that simply rephrasing toxic content with large language models can reduce toxicity while minimizing distortions in online discourse. Our findings offer critical insights and concrete strategies for policymakers and platforms to balance content distortion with harm reduction, highlighting the trade-offs of content moderation.
@@ 1 Introduction
The widespread proliferation of hateful and inflammatory content online has become an increasing concern for users, policymakers, and online platforms. As growing evidence shows that hateful online content can lead to real-life violence (Müller and Schwarz, 2021, 2022b; Bursztyn et al., 2019; Du, 2023; Cao et al., 2023), platforms have increasingly resorted to content moderation efforts to stem the tide of hateful content online. Prominent examples include the removal of Facebook accounts associated with the far-right group Proud Boys in October 2018 (e.g., NBC-News, 2018), the deletion of Alex Jones’ Twitter account in the aftermath of the Sandy Hook shooting (e.g., BBC, 2018), or most prominently, the suspension of Donald Trump’s Twitter account after the attack on the US capitol on January 6th, 2021 (e.g., Twitter, 2021; NYT, 2021). [^1] Trump’s account was only reinstated after the takeover of Twitter by Elon Musk (The Guardian, 2022). As part of the staff cuts at Twitter, Elon Musk also fired most of the content moderators on Twitter (e.g., USA-Today, 2022) Over the years, lawmakers have also started to introduce regulation of online platforms that demand the removal of Toxic online content. For example, Germany’s “Netzwerksdurchsetzungsgesetz” (BBC, 2017), the UK’s “Online Safety Bill”(e.g, Reuters, 2023b), and the “Digital Services Act” of the EU (e.g., Reuters, 2023a) mandate that online platforms are responsible for the content that is circulating on them and therefore have to take content moderation measures. As of 2020, laws that mandate removing toxic content from social media platforms had been passed in at least 25 countries (Justitia, 2020).
@@
On the one hand, concerns about hateful content and the increased demand for content moderation have motivated extensive research on automated hate speech detection (e.g., Waseem and Hovy, 2016; Hanu and Unitary team, 2020; Hartvigsen et al., 2022; Bianchi et al., 2022) and the effectiveness of content moderation efforts (e.g., Chandrasekharan et al., 2017; Jhaver et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Jiménez Durán et al., 2022; Müller and Schwarz, 2022a). One upshot of this research is that algorithms are susceptible to false positives, often triggered by swear words or otherwise innocent words that happen to often be found in the context of hate speech (e.g., Attanasio et al., 2022). On the other hand, the expansions of content moderation have been criticized as restrictions to free speech and as a distortion to online discourse (e.g., Tworek, 2021; Eidelman and Ruane, 2021; United Nations Human Rights, 2018). In particular, potentially biased applications of content rules have attracted growing criticism from politicians (e.g., Samples, 2019; Vogels et al., 2020; The Texas Tribune, 2022). Further, content moderation might also falsely target people who share encounters with racism (Lee et al., 2024). As a result, online platforms face a dilemma of seemingly contradictory objectives, which they need to balance in their content moderation efforts. This dilemma of balancing the removal of inflammatory content and the preservation of plurality of opinions is exacerbated by the fact that extensive disagreements exist about how these two objectives should be weighted.
@@
It is worth highlighting that this trade-off would persist even if the “ground truth” of hate speech was perfectly known, i.e., if there was an unbiased and universally agreed-upon measure of hate speech. [^2] In a recent paper, Solomon et al. (2024) argue that US partisan differences over what constitutes hate speech might be overstated. Even in this hypothetical scenario, content moderation would distort online content if specific topics and issues were more frequently discussed using toxic language. The trade-off only vanishes in the highly unlikely case in which the toxicity of online discourse is entirely unrelated to its content.
@@
The dilemma of content moderation is further complicated by a lack of measures to quantify the distortions to online content. While many methods exist to identify hateful content (see examples above), to the best of our knowledge, no measures exist to quantify the effect of content moderation-induced changes in online content. In this paper, we narrow this gap by putting forward and validating a measure for the distortions of online content. We formalize the idea of content distortions based on changes in the statistical moments (mean and variance) of the semantic space.
@@
As a first step, we approximate the semantic space using text embeddings based on Transformer models (Vaswani et al., 2017). These embeddings represent texts as projections into a high-dimensional Euclidean space, where their semantic similarity to other texts determines their position. Texts with similar meanings will have embeddings that are closer together than semantically unrelated texts. Embeddings have become the de-facto standard text representation in natural language processing (NLP). As of 2024, the original Vaswani et al. (2017) has received over 110,000 citations, and Transformer models have proven highly successful in countless applications (e.g., Zhu et al., 2020; Strudel et al., 2021; Han et al., 2021; Radford et al., 2023), as they have been shown to capture text semantics better than previous approaches. Transformer models also form the basis of large language models and modern machine translation. Text embeddings (as opposed to count-based methods) also have proven highly successful in computational social science (e.g., Ash and Hansen, 2023; Garg et al., 2018; Kozlowski et al., 2019; Card et al., 2022).
@@
We validate the measure’s potential on a representative sample of 5 Million US political Tweets. Our analysis proceeds in four steps. First, we show that removing toxic Tweets leads to measurable distortions in online content. Importantly, no such distortions occur if Tweets are removed at random. This result persists independent of the embedding model, toxicity score, or Twitter sample. To the best of our knowledge, our paper is the first to provide a measure to quantify this trade-off empirically. We can also show that toxicity-based content moderation distorts both the mean and the variance of the semantic space. Additionally, we find that the variance of the semantic space is reduced by increased content moderation.
@@
Second, we show that the content moderation-induced distortions are not driven by the position of toxic language in the embedding space. To this end, we build on the literature on the debiasing of text embeddings (e.g., Bolukbasi et al., 2016), and create projections of the embeddings space that are orthogonal to the toxicity scores. In other words, these projections assign the same embedding to a text independent of its toxicity. We find that our results remain virtually identical with these modified embeddings. This finding suggests that it is not the toxic language itself but rather that some topics that are disproportionally discussed in a toxic manner.
Third, we shed more light on the content moderation-induced distortions of online content using a Top2Vec topic model (Angelov, 2020). This allows us to characterize which topics are disproportionally affected by content moderation. Further, to benchmark our findings, we show that removing all Tweets with a Toxicity above a threshold of 0.8 produces distortions in online content comparable to removing 4 out of 67 topics from the data at random.
Lastly, we propose an alternative approach to content moderation that, instead of removing toxic content, rephrases Tweets in a way that strips them of their toxic language while leaving the core message intact. For this analysis, we use the language generation capabilities of large language models (LLMs) (e.g., Argyle et al., 2023). Using the rephrased Tweets, we can show that this alternative form of content moderation can measurably reduce toxicity while mitigating distortions to the semantic space.
@@
Our results break new ground by proposing a measure of content distortion that allows us to quantify the inherent trade-off in content moderation. Given the fundamental importance of this trade-off, our measure is of immediate policy relevance when it comes to the moderation of toxic and fake news content. The economic theory of multitask models (Holmstrom and Milgrom, 1991; Feltham and Xie, 1994) predicts that if principles have to choose between different objectives, only one of which is measurable, the effort will be focused on the measurable tasks. In other words, in the absence of readily available measures of content distortions, online platforms, and lawmakers are likely to put far greater emphasis on removing hateful content at the “cost” of distorting online content. Hence, our measure represents an essential puzzle piece in the debate on content moderation.
@@ 2.1 Representative US Twitter Data
For our main analysis, we use the Tweets from a representative sample of US Twitter (Siegel et al., 2021). [^4] Among others, these data were also used in Müller and Schwarz (2022a); Bose et al. (2024). The sample was created by querying the Twitter user accounts API for random numbers between 1 and $2^{32}$, the largest possible Twitter user ID at the time of collection. [^5] Twitter later switched user IDs to 64-bit. If the API returned a user account associated with the random number, the authors confirmed that the user was located in the United States. We collected the Tweets of 432,882 out of 498,901 users whose accounts were still active at the beginning of 2022. In total, this yields a dataset of ca. 400 Million Tweets. For our analysis, we removed non-English Tweets and Retweets, including those containing only links. [^6] The filtering of non-English Tweets is based on the language tags from the Perspectives API. We then finetune a BERTweet model (Nguyen et al., 2020) for the classification of political Tweets (see Appendix A for details). We use the tuned model to classify all Tweets as either political or apolitical.
@@
We create two samples for our analysis. Our main analysis is based on a sample of 5 million randomly-drawn political Tweets. [^7] As we show in our analysis, removing Tweets at random has no impact on our measure. The results therefore would be identical if we used all Tweets instead of the subsample. However, a sample of 5 Million is computationally more efficient to handle. Secondly, for robustness checks, we create a sample of 1 million randomly-drawn Tweets independent of whether they are political or apolitical. Together, the two samples provide a good approximation of political and overall Twitter content in the United States. In robustness checks we additionally consider a sample of 1 Million German and Italian Tweets from Jiménez Durán et al. (2022); Lupo et al. (2024).
@@
For our content moderation analysis, we assign three toxicity scores to each Tweet based on different models. Our baseline model is Google’s Perspectives API (Per, ). The Perspectives API has become one of the standard tools for toxicity analysis and is used by several platforms for content moderation (e.g., NYT, 2016; Forbes, 2019; Delgado, 2019). For each Tweet, the API returns six scores measuring different toxicity dimensions (toxicity, severe toxicity, identity attack, insult, profanity, and threat). The toxicity scores range from 0 (non-toxic) to 1 (highly toxic) and are roughly interpretable as the share of users who would judge a message as toxic. As is standard in the literature, we focus on Toxicity scores. [^8] As we show in our robustness tests, the results are very similar if we use any of the other toxicity dimensions. We create similar toxicity scores using the Detoxify package (Hanu and Unitary team, 2020) and the Moderation API from OpenAI (OpenAI, 2024). We show that both of these alternative models produce toxicity scores that are strongly positively correlated with our baseline measure (see Appendix Figure A.2). Additional details on the data and the toxicity measures can be found in Appendix A. Appendix Table A.3 provides examples of toxic Tweets in our data (Warning: The examples contain offensive language.)

### v2

@@ Abstract
There is an ongoing debate about how to moderate toxic speech on social media and the impact of content moderation on online discourse. This paper proposes and validates a methodology for measuring the content-moderation-induced distortions in online discourse using text embeddings from computational linguistics. Applying the method to a representative sample of 5 million US political Tweets, we find that removing toxic Tweets alters the semantic composition of content. This finding is consistent across different embedding models, toxicity metrics, and samples. Importantly, we demonstrate that these effects are not solely driven by toxic language but by the removal of topics often expressed in toxic form. We propose an alternative approach to content moderation that uses generative Large Language Models to rephrase toxic Tweets, preserving their salvageable content rather than removing them entirely. We show that this rephrasing strategy reduces toxicity while minimizing distortions in online content.
@@ 1 Introduction
The widespread proliferation of hateful and inflammatory content online has become an increasing concern for users, policymakers, and online platforms. As growing evidence shows that hateful online content can lead to real-life violence (Müller and Schwarz, 2021, 2022b; Bursztyn et al., 2019; Du, 2023; Cao et al., 2023), platforms have increasingly resorted to content moderation efforts to stem the tide of hateful content online. Prominent examples include the removal of Facebook accounts associated with the far-right group Proud Boys in October 2018 (e.g., NBC-News, 2018), the deletion of Alex Jones’ Twitter account in the aftermath of the Sandy Hook shooting (e.g., BBC, 2018), or most prominently, the suspension of Donald Trump’s Twitter account after the attack on the US capitol on January 6th, 2021 (e.g., Twitter, 2021; NYT, 2021). [^1] Trump’s account was only reinstated after the takeover of Twitter by Elon Musk (The Guardian, 2022). As part of the staff cuts at Twitter, Elon Musk also fired most of the content moderators on Twitter (e.g., USA-Today, 2022) Over the years, lawmakers have also started to introduce regulation of online platforms that regulate the removal of Toxic online content. For example, Germany’s “Netzwerksdurchsetzungsgesetz” (BBC, 2017), the UK’s “Online Safety Bill”(e.g, Reuters, 2023b), and the “Digital Services Act” of the EU (e.g., Reuters, 2023a) mandate that online platforms are responsible for the content that is circulating on them and therefore have to take content moderation measures. As of 2020, laws that mandate removing toxic content from social media platforms had been passed in at least 25 countries (Justitia, 2020).
On the one hand, concerns about hateful content and the increased demand for content moderation have motivated extensive research on automated hate speech detection (e.g., Waseem and Hovy, 2016; Hanu and Unitary team, 2020; Hartvigsen et al., 2022; Bianchi et al., 2022) and the effectiveness of content moderation efforts (e.g., Chandrasekharan et al., 2017; Jhaver et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Jiménez Durán et al., 2022; Müller and Schwarz, 2022a).
@@
On the other hand, the expansions of content moderation have been criticized as restrictions to free speech and as a distortion to online discourse (e.g., Tworek, 2021; Eidelman and Ruane, 2021; United Nations Human Rights, 2018). In particular, potentially biased applications of content rules have attracted growing criticism from politicians (e.g., Samples, 2019; Vogels et al., 2020; The Texas Tribune, 2022). It has also been shown that algorithms are susceptible to false positives, often triggered by swear words or otherwise innocent words that frequently appear in the context of hate speech (e.g., Attanasio et al., 2022). Further, content moderation might also falsely target people who share their encounters with racism (Lee et al., 2024).
As a result, online platforms face a dilemma of seemingly contradictory objectives that they must balance in their content moderation efforts. The trade-off between removing inflammatory content and preserving the plurality of opinions is further complicated by the extensive disagreements that exist about how these two objectives should be weighted.
@@
It is worth highlighting that this trade-off would persist even if the “ground truth” of hate speech was perfectly known, i.e., if there was an unbiased and universally agreed-upon measure of hate speech. [^2] In practice, hate speech detection is considered a subjective labeling task with high variation (Ross et al., 2016; Röttger et al., 2022). Even in this hypothetical scenario, content moderation would distort online content if specific topics and issues were more frequently discussed using toxic language. The trade-off only vanishes in the highly unlikely case in which the toxicity of online discourse is entirely unrelated to its content.
@@
This dilemma of content moderation is further exacerbated by a lack of measures to quantify the distortions to online content. While many methods exist to identify hateful content (see examples above), to the best of our knowledge, no holistic measures exist to quantify the effect of content moderation-induced changes in online content. In this paper, we narrow this gap by putting forward and validating a measure for the distortions of online content. We formalize the idea of content distortions based on changes in the statistical moments (mean and variance) of the semantic space.
@@
As a first step, we approximate the semantic space using text embeddings based on Transformer models (Vaswani et al., 2017), which have become the de facto standard text representation in natural language processing (NLP). As of 2025, the original Vaswani et al. (2017) has received over 200,000 citations, and Transformer models have proven highly successful in numerous applications (e.g., Zhu et al., 2020; Strudel et al., 2021; Han et al., 2021; Radford et al., 2023), as they have been shown to capture text semantics more effectively than previous approaches. Transformer models also form the basis of large language models and modern machine translation, and are capable of capturing even subtle text characteristics. Text embeddings (as opposed to count-based methods) also have proven highly successful in computational social science (e.g., Ash and Hansen, 2023; Garg et al., 2018; Kozlowski et al., 2019; Card et al., 2022). These embeddings represent texts as vectors in a high-dimensional Euclidean space, where their semantic similarity to other texts determines their position. Texts with similar meanings will have embedding vectors that are closer together than semantically unrelated texts.
@@
As we discuss in more detail later in the paper, our measure also has at least two main advantages over potential alternative measures based on cosine similarity and topic models. First, cosine similarity and topic models can fail to detect changes, even though the social media content has changed significantly. Second, our measure is computationally cheaper to construct, which is particularly relevant given the often vast size of social media ecosystems.
@@
We validate the measure’s potential on a representative sample of 5 million US political Tweets. Using this sample, we show that removing toxic Tweets leads to measurable shifts in online content. In other words, content moderation is not semantically neutral. Importantly, no such distortions occur if Tweets are removed at random. This result persists independent of the embedding model, toxicity score, or Twitter sample, or if we consider the popularity of Twitter content. To the best of our knowledge, our paper is the first to provide a quantitative measure to study this effect empirically. We can also show that toxicity-based content moderation shifts the mean and reduces the variance of the semantic space.
We benchmark the magnitude of these distortions in two ways. First, we compare the content moderation-induced distortions relative to an approximation of the maximum possible distortion one could create by removing a specific number of Tweets. We find that the removal of toxic content reaches around 20% of the maximum possible distortion. Second, we use a Top2Vec topic model (Angelov, 2020) to compare how many topics (out of 67), we would have to remove from the data to achieve comparable distortions in the semantic space as content moderation. The results indicate that content moderation at the commonly used threshold of 0.8 is comparable to removing 4 out of 67 topics from the data. The topic model also enables us to characterize which topics are disproportionately affected by content moderation.
The previously documented distortions could arise for two reasons. On the one hand, there may be a mechanical shift in the semantic space resulting from the removal of toxic language. Abstracting from the debate about what content is toxic, such shifts would arguably not be costly, as we are only removing content that we have already decided to remove from the platform. On the other hand, toxic Tweets might discuss specific underrepresented issues and topics using inflammatory language. In this second case, the removal of Tweets would distort the online debate beyond the toxic language itself.
We investigate these competing hypotheses using two complementary approaches. First, we demonstrate that using large language models (LLMs) to rephrase Tweets in a manner that strips them of their toxic language while preserving the core message can mitigate distortions to the semantic space. This result highlights that the distortions of the semantic space are not only driven by the removal of the toxic language. Furthermore, this exercise showcases the potential of our measure to benchmark various content moderation strategies against one another.
@@
Second, we build on the literature on debiasing text embeddings (e.g., Bolukbasi et al., 2016) to show that content moderation-induced distortions are not driven by the position of toxic language in the embedding space, but rather by content moderation-induced changes to online content. Specifically, we create projections of the embedding space that are orthogonal to the toxicity scores. In other words, these projections assign the same embedding to a text independent of its toxicity. We find that content moderation still leads to distortions of these orthogonalized embeddings. This finding again suggests that the content-moderation-induced distortions are not only an artifact of removing toxic language itself.
Our paper contributes to a fast-growing literature on the political effects of social media (see Zhuravskaya et al., 2020, for a review). Among others, Social media platforms have been shown to increase political polarization (Sunstein, 2017; Allcott and Gentzkow, 2017; Boxell et al., 2017; Levy, 2021; Mosquera et al., 2020), facilitate protests (Enikolopov et al., 2020; Acemoglu et al., 2017; Fergusson and Molina, 2021; Howard et al., 2011), reduce corruption and confidence in government (Enikolopov et al., 2018; Guriev et al., 2020), influence voting decisions, (Bond et al., 2012; Jones et al., 2017; Fujiwara et al., 2021), and cause offline hate crime (Müller and Schwarz, 2021, 2022b; Bursztyn et al., 2019; Cao et al., 2023).
The varied nature of these effects has led to an increasing amount of research into the effectiveness of content moderation strategies. Theoretical work by (Liu et al., 2021; Madio and Quinn, 2021; Kominers and Shapiro, 2024; Beknazar-Yuzbashev et al., 2024) as well as empirical work by (Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Müller and Schwarz, 2022a; Jiménez Durán et al., 2022) provide insights into the effects of content moderation with regard to online hate speech. A related literature has investigate interventions against online misinformation (e.g., Barrera et al., 2020; Henry et al., 2022, 2023)
However, none of the above studies investigates the potentially adverse consequences of such interventions on online expression. So far, these questions have been tackled using surveys to document the popular support or agreement on which content should be removed (Kozyreva et al., 2023; Solomon et al., 2024; Munzert et al., 2025). While important, the popular agreement on specific content moderation measures can not be used to judge the cost of content moderation. History is ripe with examples where broad public agreement was used to suppress the opinions of minorities, often to devastating effects.
@@
Our results break new ground by proposing a measure of content distortion that enables benchmarking different content moderation approaches and allows us to quantify the inherent trade-off in content moderation. Given the fundamental importance of this trade-off, our measure is of immediate policy relevance when it comes to the moderation of toxic and fake news content. The economic theory of multitask models (Holmstrom and Milgrom, 1991; Feltham and Xie, 1994) predicts that if principals must choose between different objectives, only one of which is measurable, effort will be focused on the measurable tasks. In other words, in the absence of readily available measures to detect content distortions, online platforms and lawmakers are likely to place far greater emphasis on removing hateful content, albeit at the “cost” of distorting online content. Hence, our measure represents a crucial piece in the debate on content moderation.
@@ 2.1 Representative US Twitter Data
For our main analysis, we use the Tweets from a representative sample of US Twitter (Siegel et al., 2021). [^4] Among others, these data were also used in Müller and Schwarz (2022a) and Bose et al. (2024). The sample was created by querying the Twitter user accounts API for random numbers between 1 and $2^{32}$, the largest possible Twitter user ID at the time of collection. [^5] Twitter later switched user IDs to 64-bit. If the API returned a user account associated with the random number, the authors confirmed that the user was located in the United States. We collected the Tweets of 432,882 out of 498,901 users whose accounts were still active at the beginning of 2022. In total, this yields a dataset of ca. 400 Million Tweets. For our analysis, we removed non-English Tweets and Retweets, including those containing only links. [^6] The filtering of non-English Tweets is based on the language tags from the Perspectives API. We then finetune a BERTweet model (Nguyen et al., 2020) for the classification of political Tweets and classify all Tweets as either political or apolitical (see Appendix A for details).
@@
We create two samples for our analysis. Our main analysis is based on a sample of 5 million randomly-drawn political Tweets. [^7] As we show in our analysis, removing Tweets at random has no impact on our measure. The results, therefore, would be identical if we used all Tweets instead of the subsample. However, a sample of 5 million is computationally more efficient to handle. Secondly, for robustness checks, we create a sample of 1 million randomly-drawn Tweets independent of whether they are political or apolitical. Together, the two samples provide a good approximation of political and overall Twitter content in the United States. In robustness checks, we additionally consider a sample of 1 million German and Italian Tweets from Jiménez Durán et al. (2022) and Lupo et al. (2024).
@@
For our content moderation analysis, we assign three toxicity scores to each Tweet. Our baseline model is Google’s Perspective API (Perspective, 2017). The Perspectives API has become one of the standard tools for toxicity analysis and is used by several platforms for content moderation (e.g., NYT, 2016; Forbes, 2019; Delgado, 2019). For each Tweet, the API returns six scores measuring different toxicity dimensions (toxicity, severe toxicity, identity attack, insult, profanity, and threat). The toxicity scores range from 0 (non-toxic) to 1 (highly toxic) and are roughly interpretable as the share of users who would judge a message as toxic. As is standard in the literature, we focus on Toxicity scores. [^8] As we show in our robustness tests, the results are very similar if we use any of the other toxicity dimensions. We create similar toxicity scores using the Detoxify package (Hanu and Unitary team, 2020) and the Moderation API from OpenAI (OpenAI, 2024). We show that both alternative models produce toxicity scores that are strongly positively correlated with our baseline measure (see Appendix Figure A.2). Additional details on the data and the toxicity measures can be found in Appendix A. Table 1 provides examples of toxic Tweets in our data (Warning: The examples contain offensive language.)
Table 1: Examples Toxic Tweets
The Top2Vec algorithm proceeds in three steps. First, the texts are transformed into embedding vectors based on pre-trained models. In our particular case, we make use of the universal-sentence-encoder model (Cer et al., 2018). Second, the embeddings are projected into a lower-dimensional space using UMAP (McInnes et al., 2018). This helps to overcome the sparsity of the higher-dimensional space. Third, the reduced embeddings are clustered into topics using HDBSCAN (McInnes et al., 2017), which identifies the centroids for each topic vector. We specify that HDBSCAN should only create clusters with at least 1500 observations. Top2Vec creates a unique assignment between texts and topics. The topic words are then derived based on the word vectors that are closest to the topic centroid. We report the most important topic words as well as the assigned topic labels in Appendix Table A.3.